The length of the rectangle is 11 less than twice its width. If the perimeter is 86 what is the length and width of the rectangle?
Answer by rm29924(97)

Let X = the width
the length is 11 less than twice the width so
Length = 2x-11
the perimeter of a rectangle = 2W + 2L so
2X + 2(2X-11) = 86
2X + 4X -22 = 86
add 22 to both sides
2X + 4X = 108
6X = 108
divide both sides by 6
X = 18
check the answer
2(18) + 2(2(18)-11) = 86
36 + 2(36-11) = 86
36 + 2(25) = 86
36 + 50 = 86
86 = 86